kaly:

Stacey currently has an account balance of $1,007.78. Her initial deposit on the account was $616 and it earned 5.3% simple interest. How long has Stacey held the account?

Shadow:

Hello

Shadow:

Let me know if you are here and I can help you solve this problem real quick.

kaly:

Yes I am here. I have been trying to figure this answer out for almost an hour and cant figure the answer out!!!

Shadow:

Well, it is quite simple. You will be using the simple interest formula: \[A = P(1 + rt)\] A = Total Accrued Amount (principal + interest) P = Principal Amount I = Interest Amount r = Rate of Interest per year in decimal; r = R/100 R = Rate of Interest per year as a percent; R = r * 100 t = Time Period involved in months or years

Shadow:

Does that look familiar to you?

kaly:

Parts of that does look very familiar. Not all of it but alot yes.

Shadow:

Let me break it down for you. P, or the principal amount, is the $ you started with, which is $616 in this case. A, or the accrued amount, is the $ you started + the $ you earned (also known as I, the interest amount). r is simply the rate given that shows you how much money you get. Since this problem doesn't state a specific period such as daily, or monthly, we will assume that it is the standard annually (aka yearly, or per year). t is what we are trying to solve for, which is the amount of time, in years, that Stacey had the account.

Shadow:

Do you see how we might put this problem together?

kaly:

Yes I do and I do believe that I can figure out the answer now myself.. Thank you for helping!!